

********************************************************************
UARS WEEKLY STATUS REPORT - 25 AUG 2000
********************************************************************
***********
SOLSTICE
***********
SOLSTICE is in good health, collecting solar data on all available
orbital tracking periods and stellar calibration observations whenever
possible. Our solar data is not affected by the larger than normal
pointing errors since they are within our calibrations. Additionally, we
continue to gather enough data to meet all the SOLSTICE science
objectives during this period of reduced tape recorder operations.
Level2 processing is halted due to bugs in the OA_SSPP_ATT service. CDHF
and CSC personnel are working on the problem.
***********
SUSIM
***********
SUSIM continues to gather a normal set of daily solar scans
scheduled for periods of predicted telemetry.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
The PEM zenith spectrometers are activate in daylight mode.
At the present time, PEM energy deposition is not being generated.
Recent data configuration changes have affected PEM software causing multiple
failures. These failures are being reviewed by the PEM team. In order to
correct for new telemetry requirements, PEM will need to update processing
code with a patch delivery. The delivery patch will not include changes which
would allow the PEM software to execute on a DEC Alpha machine.
Updated HEPS calibration parameters have been finalized and are ready for
delivery. Updated MEPS quality calibration files have been finalized and are
ready for delivery. Level 3 software errors, discovered by Dr. Dean Pesnell,
have been fixed; his suggestions have been incorporated with in the PEM
production code. Both updated HEPS and MEPS calibration data, along with the
Level 3 code changes, will influence the in-situ energy deposition predictions,
so new versions of PEM energy deposition Level 3 files will be generated.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG participated in the meeting on Wednesday, 23 August,
held to discuss the status of the ongoing efforts by the
Flight Dynamics Facility and the Flight Operations Team to
improve calibration of UARS Fixed Head Star Tracker 1.
Processing of instrument Daily Activity Plans continues with no
problems. We are continuing to coordinate with the Flight
Operations Team and to assist the Project Scientist in the
scheduling of instrument operations.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 12 August 2000 (Orbit 48752, GMT Day 225) through 18 August
2000 (Orbit 48855, GMT Day 231).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
ACRIM 226/1433 48775 Off
SOLSTICE 226/1433 48775 Off
SUSIM 226/1433 48775 Off
ACRIM 229/0027 48811 On
SOLSTICE 229/0027 48811 On
SUSIM 229/1451 48820 On
PEM 230/1129 48833 On - Day Only
The beta angle ranged between the angles of 13.4 degrees to 15.5 degrees for
this report period. The beta angle is now increasing toward a maximum angle of
+49 degrees on 31 August 2000 (DOY 244).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 1.0 to 1.3 degrees C. Additional battery performance data for each
battery on 11 August 2000 and 18 August 2000 are provided in Appendix D.
The clock error ranged from -3.2 to 13.9 msec during this report period. Clock
rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
¥ OUT-OF-LIMITS = 00
¥ ALERTS = 00
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
¥ FOT = 17
¥ SUSIM = 02
¥ PEM = 02
¥ ACRIM = 01
¥ SOLSTICE = 01
¥ SSPP = 01
¥ HRDI = 01
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
¥ FOT = 01
Two (2) Anomaly Reports remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
· Transponder A: 00/153 - 2287.496961
· Transponder B: 00/213 - 2287.499288
GROUND SYSTEM OPERATION
One (1) operational support problem occurred this period. Anomaly Reports are
included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 00/230 show the observatory data loss to be 297 hours, 52
minutes, 39.185 seconds (an increase of 07 minutes 50.208 seconds since the
last report period). This is a 0.40 percent data loss which equals a 99.60
percent data capture for the mission.
The increase of 07 minutes 50.208 seconds occurred over three (3) days:
· 02 minutes 51.296 seconds on Day 226 due to no acquisition.
· 03 minutes 12.768 seconds on Day 228 due to no acquisition.
· 01 minutes 46.144 seconds on Day 229 due to no acquisition.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT continues to operate FHST-1 for approximately one orbit each day. By
cycling the tracker on and off for such a short duration the degradation seems
to have been reduced significantly. Scale factor calibrations are being
provided on a regular basis from FDF and are being uplinked when received. It
appears that the FOT has been able to stabilize the performance of FHST-1 for
now and continues to work on establishing another viable attitude control
solution should the tracker fail.
MEETINGS
The FOT attended a meeting to discuss FHST-1 on 14 and 16 August 2000.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for September 17, 2000 (DOY 261). This
maneuver will be a Reverse to Forward Flight maneuver (Southern Hemisphere
viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 18 Aug 2000
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
48761 225/1529 00-404 FHST-1 Power ON
48762 225/1750 00-404 FHST-1 Power OFF
48772 226/0848 00-412 HRDI STANDBY
48772 226/0928 00-413 PEM Power OFF
48775 226/1433 00-414 SSPP Power OFFSOLSTICE,SUSIM,ACRIM
48776 226/1532 00-404 FHST-1 Power ON
48777 226/1753 00-404 FHST-1 Power OFF
48789 227/1210 00-415 FHST1 Power ON
48789 227/1252 00-416 MACS YAW MANUEVER FWD TO REV
48792 22/1628 00-415 FHST1 Power OFF
48794 227/2102 00-417 FHST1 CALIBRATION
48806 228/1550 00-388 FHST-1 Power ON
48809 228/2011 00-388 FHST-1 Power Off
48811 229/0027 00-418 ACRIM Power ON
48811 229/0027 00-419 SOLSTICE Power ON
48820 229/1451 00-420 SUSIM Power ON
48832 230/0948 00-420 SUSIM Power ON(COMPLETE)
48833 230/1129 00-359 PEM Power ON(dayonly)
48835 230/1452 00-421 FHST-1 Scale factor update
38836 230/1557 00-422 FHST-1 Power ON
48837 230/1744 00-422 FHST-1 Power OFF
48853 231/1827 00-423 NBTR NOOPS D-232
48854 231/2010 00-424 FHST-1 FHST CALIBRATION
48854 231/2010 00-422 FHST-1 Power ON
48855 231/2154 00-422 FHST-1 Power OFF
_______________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 18 Aug 2000
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
None this report period
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None this report period
_______________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 18 Aug 2000
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
None this report period
___________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 18 Aug 2000
11 Aug 2000 - GMT Day 224 Beta = 13.4 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN
EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- --------- ---------- ------- ------- ------
1 +459.2/+168.0 -3.05/-4.62 +0.0/+0.0 15.2 0.000 0.0
2 +33.6/-44.8 +5.79/+3.87 +29.6/-14.4 24.3 1.040 19.6
3 +22.4/-33.6 +4.83/+2.60 +28.8/-15.2 24.3 1.037 19.2
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---------- ----------
no changes V/T 5
_______________________________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 18 Aug 2000
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-040 FOT 48854 231/2111 22831 DATA DROPOUT
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-037 FOT see list Aug 00 14170 Generic Late Acq.
00-032 FHST 48211 188/2242 N/A FHST 2 Anomaly
SEVNTFW00/32